With summer approaching, the number of air conditioners sold is expected to be double that of stoves and refrigerators combined. Stoves sell for $350, refrigerators for $425, and air conditioners for $525, and sales of $14,375 are expected. If stoves and refrigerators sell in equal numbers, how many of each appliance should be stocked?
Answer by stanbon(75887)

With summer approaching, the number of air conditioners sold is expected to be double that of stoves and refrigerators combined. Stoves sell for $350, refrigerators for $425, and air conditioners for $525, and sales of $14,375 are expected.
Quantity EQuation: A = 2(S + R)
----
If stoves and refrigerators sell in equal numbers, how many of each appliance should be stocked?
Quantity Eqution: S = R
---
Substitute to get A = 2(2S) = 4S
Value Equation: 350S + 425S+ 525(4S) = 14,375
----
2875S = 14375
--------------------
S = 5 (# of stoves to stock)
R = S = 5 (# of refrigerators to stock)
A = 4S = 20 (# of air conditioners to stock)
===============
Cheers,
Stan H.
===============